OpenStudy (itsbribro):

Points F and G lie on the graph of the equation y = -0.25x. What are the y-coordinates for each point? Point F: (4,____) Point G: (-8,____)

OpenStudy (anonymous):

substitute in your equation by x=4 to get y-coor. for point F and do the same for point G

OpenStudy (itsbribro):

so i would \[y=-0.25x\] would change into \[-8=-0.25(4)\] Am i right?

OpenStudy (anonymous):

y=-.25x when x=4 then y=-.25*4=-1 -1 is the y-coor. for this point F

OpenStudy (itsbribro):

so -1 would be y's missing number?

OpenStudy (itsbribro):

or g's missing number

OpenStudy (anonymous):

it is the y's missing no for point F

OpenStudy (itsbribro):

can you help me out with G then please?

OpenStudy (itsbribro):

@esamalaa

OpenStudy (itsbribro):

i would just take .25 *-8 and get my answer?

OpenStudy (itsbribro):

which is -2